# Persistent Homology for High-dimensional Data Based on Spectral Methods

### v2

@@ 7.1 Synthetic benchmark
Figure 6: Loop detection score for selected methods on synthetic datasets in ambient $\mathbb{R}^{50}$. More experimental results can be found in Figures S19 – S28. Recommended methods in bold.
@@
On the circle dataset in $\mathbb{R}^{50}$, persistent homology with all distance metrics found the correct hole when the noise level $\sigma$ was very low (Figure 5). However, as the amount of noise increased, the performance of Euclidean distance quickly deteriorated, reaching zero score at $\sigma\approx 0.2$. Most other distances outperformed the Euclidean distance, at least in the low noise regime. Fermat distance did not have any effect, and neither did DTM distance, which collapsed at $\sigma\approx 0.15$ due to our thresholding (Figure 5a). Geodesics, UMAP/$t$-SNE graph, and core distance offered only a modest improvement over Euclidean (Figure 5b) highlighting that many $k$NN-graph-based distances cannot handle high-dimensional noise. In contrast, embedding-based distances performed very well on the circle (Figure 5c), but have obvious limitations: for example, a 2D embedding cannot possibly have a void. UMAP with higher embedding dimension struggled with loop detection on surfaces and the torus’ void (Appendix L). Finally, all spectral methods (effective resistance, diffusion, and Laplacian Eigenmaps) showed similarly excellent performance (Figure 5d).
@@ 7.2 Detecting cycles in single-cell data
The Malaria dataset is expected to contain two cycles: the parasite replication cycle in red blood cells, and the parasite transmission cycle between human and mosquito hosts. Following Howick et al. [37], we based all computations for this dataset (and all derived distances) on the correlation distance instead of the Euclidean distance. Persistent homology based on the correlation distance itself failed to correctly identify the two ground-truth cycles and DTM produced representatives that only roughly approximate the two ground truth cycles (Figure 8a,b). But both effective resistance and diffusion distance successfully uncovered both cycles, with $s_{2}>0.9$ (Figure 8c,d).
@@
8 Discussion
In this work we asked how to use persistent homology on high-dimensional noisy datasets which are very common in real-world applications even if the intrinsic data dimensionality is low. We found spectral methods to be the optimal approach. We demonstrated that, as the dimensionality of the data increases, the main problem for persistent homology shifts from handling outliers to handling noise dimensions (Section 4). We used a synthetic benchmark to show that traditional persistent homology and many of its existing extensions struggle to find the correct topology in this setting. Our main finding is that spectral methods based on the $k$NN graph, such as the effective resistance and diffusion distances, still work well (Section 7.1). Furthermore, we view it as an advantage that we found existing methods that are able to handle the important problem of high-dimensional noise. We derived an expression for effective resistance based on the eigendecomposition of the graph Laplacian, and demonstrated that it combines all but the most local diffusion distances (Section 6). As a result, effective resistance only needs a single hyperparamter, the number of neighbors. For this reason and due to its greater outlier resistance (Appendix K), we tend to recommend effective resistance over diffusion distances. Finally, we showed that spectral distances outperform all competitors on single-cell data (Section 7.2).
@@ 8 Discussion
In the real-world applications, it was important to look at representatives of detected holes as some holes were persistent, but arguably incorrect. That said, each hole homology class has many different representative cycles, making interpretation difficult. Given ground-truth cycles, an automatic procedure for evaluating cycle correctness remains an interesting research question.
@@
Persistent homology can only detect topology, which is often a useful global level of abstraction. However, it may therefore fail to distinguish some non-isomorphic point clouds [64].
@@
Dimensionality reduction methods are designed to handle high-dimensional data. $t$-SNE and UMAP indeed performed well on many datasets, but on average worse than spectral distances on the real data. Moreover, UMAP struggled with the surface of the 3D toy datasets and the torus’ void (Appendix L). Finally, they require the choice of an embedding dimension and are known to produce artifacts [11, 74], e.g., leading to poor scores in the noiseless setting in Figures S13S18. In contrast, spectral distances on the $\text{s}k\text{NN}$ graph worked well without a low-dimensional embedding (Section 7).
@@
Using effective resistance or diffusion distances is easy in practice as their computation time is dwarfed by that of the persistent homology (Table S4), which scales as $\mathcal{O}(n^{3(\delta+1)})$ for $n$ points and topological holes of dimension $\delta$ [51]. This high complexity of persistent homology aggravates other problems of high-dimensional datasets as dense sampling in high-dimensional space would require a prohibitively large sample size. Combining persistent homology with non-Euclidean distance measures could mitigate this problem via the approach of Bendich et al. [6], who performed subsampling after computation of the distance matrix. This is a particularly attractive avenue for future research.
@@
We thank Enrique Fita Sanmartin, Facundo Mémoli, and Ulrike von Luxburg for productive discussions on effective resistance and persistent homology, and Benjamin Dunn, Erik Hermansen, and David Klindt for helpful discussions on combining persistent homology with other dissimilarities than Euclidean distance. We thank Bastian Rieck for feedback on the manuscript. Moreover, we thank Sten Linnarsson and Miri Danan Gotthold for sharing the scVI representation of their pallium data.
This work was funded by Deutsche Forschungsgemeinschft (DFG, German Research Foundation) via Germany’s Excellence Strategy (Excellence cluster 2064 “Machine Learning — New Perspectives for Science”, EXC 390727645; Excellence cluster 2181 “STRUCTURES”, EXC 390900948), the German Ministry of Science and Education (BMBF) via the Tübingen AI Center (01IS18039A), the Gemeinnützige Hertie-Stiftung, and the National Institutes of Health (UM1MH130981). The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
@@ Appendix B Noise in high-dimensional spaces eventually dominates structure
In this section, we formally show that the Euclidean distances and thus the traditional persistence diagrams eventually get dominated by noise as the number of ambient dimensions grows. The main results are Corollary B.5 and Corollary B.7.
We being with some helpful lemmata.
@@
Each entry of $\varepsilon_{1,d}-\varepsilon_{2,d}$ is identically, independently, normally distributed with zero mean and variance $2\sigma^{2}$. Moreover, the squared Euclidean distance is the sum of the squared entries of the vector. This implies the first statement. By the strong law of large numbers $\|\varepsilon_{1,d}-\varepsilon_{2,d}\|^{2}/d$ converges almost surely to $2\sigma^{2}$. By the continuous mapping theorem, this implies almost sure convergence of $\|\varepsilon_{1,d}-\varepsilon_{2,d}\|/\sqrt{d}$ to $\sqrt{2}\sigma$. ∎
@@
Let $\{X_{d}\}$ be a sequence of random variables whose means and variances exist. If $\mathbb{E}(X_{d})$ converges to a constant $c$ and $\textnormal{Var}(X_{d})\to 0$ as $d\to\infty$, then $\{X_{d}\}$ converges to $c$ in squared mean, i.e., $\mathbb{E}\big{(}(X_{d}-c)^{2}\big{)}\to 0$.
@@
$\mathbb{P}(|X_{d}-X|>\varepsilon)=\mathbb{P}((X_{d}-X)^{2}>\varepsilon^{2})\leq\mathbb{E}((X_{d}-X)^{2})/\varepsilon^{2}\to 0$
@@
Now we show that for sufficiently many noise dimensions the distance between any two points gets dominated by the noise. This follows from the Pythagorean theorem, because most noise dimensions are orthogonal to the difference of the noise-free points.
@@
Let $x_{1}$ and $x_{2}$ be two points in $\mathbb{R}^{d^{\prime}}$. Let $\{\iota_{d}\}$ be a sequence of isometries $\iota_{d}:\mathbb{R}^{d^{\prime}}\to\mathbb{R}^{d}$ for $d\geq d^{\prime}$. Let $\{\varepsilon_{1,d}\}$ and $\{\varepsilon_{2,d}\}$ be two sequences of $d$-dimensional mutlivariate normally distributed random variables $\varepsilon_{i,d}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I}_{d})$ starting at $d=d^{\prime}$. Let $\{\Delta_{d}\}$ be the sequence of Euclidean distances between $\iota_{d}(x_{1})+\varepsilon_{1,d}$ and $\iota_{d}(x_{2})+\varepsilon_{2,d}$. Then $\{\Delta_{d}^{2}/d\}$ converges to $2\sigma^{2}$ in squared mean.
@@
$\Delta_{d}=\|\iota_{d}(x_{1})+\varepsilon_{1,d}-(\iota_{d}(x_{2})+\varepsilon_{2,d})\|=\|(\iota_{d}(x_{1})+\varepsilon_{1,d}-\varepsilon_{2,d})-\iota_{d}(x_{2})\|.$
@@
Let $\alpha$ be the angle that $\varepsilon_{d}$ makes with the vector $x_{2}-x_{1}$. Then the part of $\varepsilon_{d}$ normal to $x_{2}-x_{1}$ is of length $n=\sin(\alpha)l$ and the part parallel to it is of (signed) length $p=\cos(\alpha)l$. By the Pythagorean theorem, we have
$\Delta_{d}^{2}$ $=n^{2}+(\delta-p)^{2}=n^{2}+\delta^{2}-2\delta p+p^{2}=l^{2}(\sin(\alpha)^{2}+\cos(\alpha)^{2})+\delta^{2}-2\delta\cos(\alpha)l$
$=l^{2}+\delta^{2}-2\delta\cos(\alpha)l.$ (9)
Since $\varepsilon_{d}$ is isometrically distributed, we can assume that $x_{2}-x_{1}$ is parallel to $(1,0,\dots,0)$ without loss of generality. Then $\cos(\alpha)l=\varepsilon^{\prime}$, where $\varepsilon^{\prime}\sim\mathcal{N}(0,2\sigma^{2})$ is the first entry of $\varepsilon_{d}$.
Thus, $\mathbb{E}(\Delta_{d}^{2})=2\sigma^{2}d+\delta^{2}$ by Lemma B.1and $\mathbb{E}(\Delta_{d}^{2}/d)\to 2\sigma^{2}$ for $d\to\infty$. For the variance, we obtain
@@
These statements show that noise will eventually dominate the Euclidean distances and drive traditional persistent homology if the ambient dimension is large enough. In Figure 7 we saw empirically that tens of ambient dimensions already severely impact traditional persistent homology, while spectral distance offer more noise robustness. This also holds true in hundreds to thousands of ambient dimensions. Compared to the Euclidean distance, spectral distance can detect the correct topology in more than an order of magnitude higher dimensionality, before they eventually fail, too (Figure S1).

### v3

@@ 7.1 Synthetic benchmark
Figure 7: Loop detection score for selected methods on synthetic datasets in ambient $\mathbb{R}^{50}$. More experimental results can be found in Figures S23 – S33. Recommended methods in bold.
@@
On the circle dataset in $\mathbb{R}^{50}$, persistent homology with all distance metrics found the correct hole when the noise level $\sigma$ was very low (Figure 6). However, as the amount of noise increased, the performance of Euclidean distance quickly deteriorated, reaching zero score at $\sigma\approx 0.2$. Most other distances outperformed the Euclidean distance, at least in the low noise regime. Fermat distance did not have any effect, and neither did DTM distance, which collapsed at $\sigma\approx 0.15$ due to our thresholding (Figure 6a). Geodesics, UMAP/$t$-SNE graph, and core distance offered only a modest improvement over Euclidean (Figure 6b) highlighting that many $k$NN-graph-based distances cannot handle high-dimensional noise. In contrast, embedding-based distances performed very well on the circle (Figure 6c), but have obvious limitations: for example, a 2D embedding cannot possibly have a void. UMAP with higher embedding dimension struggled with loop detection on surfaces and the torus’ void (Appendix O). Finally, all spectral methods (effective resistance, diffusion, and Laplacian Eigenmaps) showed similarly excellent performance (Figure 6d).
@@ 7.2 Detecting cycles in single-cell data
The Malaria dataset is expected to contain two cycles: the parasite replication cycle in red blood cells, and the parasite transmission cycle between human and mosquito hosts. Following Howick et al. [43], we based all computations for this dataset (and all derived distances) on the correlation distance instead of the Euclidean distance. Persistent homology based on the correlation distance itself failed to correctly identify the two ground-truth cycles and DTM produced representatives that only roughly approximate the two ground truth cycles (Figure 9a,b). Both effective resistance and diffusion distance successfully uncovered both cycles with $s_{2}>0.9$ (Figure 9c,d).
@@
8 Limitations and future work
@@
In the real-world applications, it was important to look at representatives of detected holes as some holes were persistent, but arguably incorrect. That said, each homology class has many different representative cycles, making interpretation difficult. Given ground-truth cycles, an automatic procedure for evaluating cycle correctness remains an interesting research question.
@@
Persistent homology can only detect topology, which is often a useful global level of abstraction. However, it may therefore fail to distinguish some non-isomorphic point clouds [75]. There exist dedicated measures for detecting isometry [7, 87, 47].
@@
Dimensionality reduction methods are designed to handle high-dimensional data. $t$-SNE and UMAP indeed performed well on many datasets, but on average worse than spectral distances on the real data. Moreover, UMAP struggled with the surface of the 3D toy datasets and the torus’ void (Appendix O). Finally, they require the choice of an embedding dimension and are known to produce artifacts [49, 12, 85], e.g., leading to poor scores in the noiseless setting in Figures S17S22. In contrast, spectral distances on the symmetric $k$NN graph worked well without a low-dimensional embedding (Section 7).
@@
Using effective resistance or diffusion distances is easy in practice as their computation time $O(n^{3}$) is dwarfed by that of the persistent homology (Table S5), which scales as $\mathcal{O}(n^{3(\delta+1)})$ for $n$ points and topological holes of dimension $\delta$ [59]. This high complexity of persistent homology aggravates other problems of high-dimensional datasets as dense sampling in high-dimensional space would require a prohibitively large sample size (recall that spectral methods needed a high sampling density for good performance on some of our datasets such as the torus). Combining persistent homology with non-Euclidean distance measures could mitigate this problem via the approach of Bendich et al. [6], who performed subsampling after computation of the distance matrix. This is a particularly attractive avenue for future research.
Both effective resistance and diffusion distances require the choice of hyperparameters. However, effective resistance only needs a single hyperparameter: the number of $k$NN neighbors. For this reason and due to its greater outlier resistance (Appendix N), we tend to recommend effective resistance over diffusion distances, but a principled criterion when to use which of the two is still missing.
Moreover, we do not have a theoretical proof that spectral distances mitigate the curse of dimensionality. Such a proof may be achieved in the future taking inspiration from the stability results in [11, 42, 77] and, more generally, spectral perturbation theory.
Our empirical results focus on benchmarking which distances identify the correct topology in the presence of high-dimensional noise. Therefore, we only considered datasets with known ground-truth topology. The next step will be to use spectral distances to detect non-trivial topology in real-world exploratory contexts.
High-dimensional data and thus application areas for our improved topology detection pipeline are becoming ubiquitous. Within biology, we see possible applications for our method in other single-cell omics modalities, population genomics, or neural activity data [31, 40]. Beyond biology, we believe that our approach can improve the topological analysis of artificial neural network activations [60], and in general be used to detect topology of any high-dimensional data, e.g. in the climate sciences, in astronomical measurements, or wearable sensor data.
9 Conclusion
In this work we asked how to use persistent homology on high-dimensional noisy datasets which are very common in real-world applications even if the intrinsic data dimensionality is low. We found spectral methods to be the optimal approach. We demonstrated that, as the dimensionality of the data increases, the main problem for persistent homology shifts from handling outliers to handling noise dimensions (Section 4). We used a synthetic benchmark to show that traditional persistent homology and many of its existing extensions struggle to find the correct topology in this setting. Our main finding is that spectral methods based on the $k$NN graph, such as the effective resistance and diffusion distances, still work well (Section 7.1). Furthermore, we view it as an advantage that we found existing methods that are able to handle the important problem of high-dimensional noise. We derived an expression for effective resistance based on the eigendecomposition of the graph Laplacian, and demonstrated that it combines all but the most local diffusion distances (Section 6). Finally, we showed that spectral distances outperform all competitors on single-cell data (Section 7.2).
@@ 9 Conclusion
We thank Enrique Fita Sanmartin and Ulrike von Luxburg for productive discussions on effective resistance and persistent homology, and Benjamin Dunn, Erik Hermansen, and David Klindt for helpful discussions on combining persistent homology with other dissimilarities than Euclidean distance. Moreover, we thank Sten Linnarsson and Miri Danan Gotthold for sharing the scVI representation of their pallium data. Final thanks go to Bastian Rieck for feedback on the writing.
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via Germany’s Excellence Strategy (Excellence cluster 2064 “Machine Learning — New Perspectives for Science”, EXC 390727645; Excellence cluster 2181 “STRUCTURES”, EXC 390900948), the German Ministry of Science and Education (BMBF) via the Tübingen AI Center (01IS18039A), the Gemeinnützige Hertie-Stiftung, and the National Institutes of Health (UM1MH130981). The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
@@ Appendix B Noise in high-dimensional spaces eventually dominates structure
In this section, we rigorously prove the intuitive result that the Euclidean distances — and thus the traditional persistence diagrams — eventually get dominated by noise as the number of ambient dimensions grows (assuming homoscedastic noise). The main results are Corollary B.5 and Corollary B.7. Our analysis is similar to the concurrent work of Hiraoka et al. [42]. Their analysis extends ours in that they consider an arbitrary noise distribution, the Čech complex, and make more precise statements on how persistent homology gets impaired by high-dimensional noise.
We begin with some helpful lemmata.
@@
All entries of $\varepsilon_{1,d}-\varepsilon_{2,d}$ are i.i.d. Gaussian random variables with zero mean and variance $2\sigma^{2}$. The squared Euclidean distance is the sum of the squared entries of the vector, which implies the first statement. By the strong law of large numbers, $\|\varepsilon_{1,d}-\varepsilon_{2,d}\|^{2}/d$ converges almost surely to $2\sigma^{2}$. By the continuous mapping theorem, this implies almost sure convergence of $\|\varepsilon_{1,d}-\varepsilon_{2,d}\|/\sqrt{d}$ to $\sqrt{2}\sigma$. ∎
@@
Let $\{X_{d}\}$ be a sequence of random variables with finite means and variances. If $\mathbb{E}(X_{d})$ converges to a constant $c$ and $\textnormal{Var}(X_{d})\to 0$ as $d\to\infty$, then $\{X_{d}\}$ converges to $c$ in squared mean, i.e., $\mathbb{E}\big{(}(X_{d}-c)^{2}\big{)}\to 0$.
@@
$\mathbb{P}\big{(}|X_{d}-X|>\varepsilon\big{)}=\mathbb{P}\big{(}(X_{d}-X)^{2}>\varepsilon^{2}\big{)}\leq\mathbb{E}\big{(}(X_{d}-X)^{2}\big{)}/\varepsilon^{2}\to 0$
@@
Now we show that for sufficiently many noise dimensions the distance between any two points gets dominated by the noise. This follows from the Pythagorean theorem, because most noise dimensions are orthogonal to the difference of the noise-free points. A similar result can be found in Hall et al. [38].
@@
Let $x_{1}$ and $x_{2}$ be two points in $\mathbb{R}^{d^{\prime}}$. Let $\{\iota_{d}\}$ be a sequence of isometries $\iota_{d}:\mathbb{R}^{d^{\prime}}\to\mathbb{R}^{d}$ for $d\geq d^{\prime}$. Let $\{\varepsilon_{1,d}\}$ and $\{\varepsilon_{2,d}\}$ be two sequences of $d$-dimensional multivariate normally distributed random variables $\varepsilon_{i,d}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I}_{d})$ starting at $d=d^{\prime}$. Let $\{\Delta_{d}\}$ be the sequence of Euclidean distances between $\iota_{d}(x_{1})+\varepsilon_{1,d}$ and $\iota_{d}(x_{2})+\varepsilon_{2,d}$. Then $\{\Delta_{d}^{2}/d\}$ converges to $2\sigma^{2}$ in squared mean.
@@
$\Delta_{d}=\|\iota_{d}(x_{1})+\varepsilon_{1,d}-\big{(}\iota_{d}(x_{2})+\varepsilon_{2,d}\big{)}\|=\|\big{(}\iota_{d}(x_{1})+\varepsilon_{1,d}-\varepsilon_{2,d}\big{)}-\iota_{d}(x_{2})\|.$
@@
Let $\alpha$ be the angle that $\varepsilon_{d}$ makes with the vector $x_{2}-x_{1}$. Then, by the law of cosines,
$\Delta_{d}^{2}=l^{2}+\delta^{2}-2\delta\cos(\alpha)l.$ (9)
Since $\varepsilon_{d}$ is isometrically distributed, we can assume that $x_{2}-x_{1}$ is parallel to $(1,0,\dots,0)$ without loss of generality. Then $\cos(\alpha)l=\varepsilon^{\prime}$, where $\varepsilon^{\prime}\sim\mathcal{N}(0,2\sigma^{2})$ is the first entry of $\varepsilon_{d}$. Thus, $\mathbb{E}(\Delta_{d}^{2})=2\sigma^{2}d+\delta^{2}$ by Lemma B.1 and $\mathbb{E}(\Delta_{d}^{2}/d)\to 2\sigma^{2}$ for $d\to\infty$.
For the variance, we obtain
@@
These statements show that noise will eventually dominate the Euclidean distances and drive traditional persistent homology if the ambient dimension is large enough. In Figure 8 we saw empirically that tens of ambient dimensions already severely impact traditional persistent homology, while spectral distance offer more noise robustness. This also holds true in hundreds to thousands of ambient dimensions. Compared to the Euclidean distance, spectral distance can detect the correct topology in more than an order of magnitude higher dimensionality, before they eventually fail too (Figure S1).